40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jan 9 09:18:47 2017, Steve, Update, PEM, power glitch MondayMorning.png
    Reply  Tue Jan 10 16:20:11 2017, Steve, Update, CDS, power glitch......Raid is rebuilding rebuilding_in_progress.png
    Reply  Tue Jan 10 21:47:00 2017, rana, Update, CDS, power glitch 
       Reply  Tue Jan 10 22:55:43 2017, gautam, Update, CDS, power glitch - recovery steps 
          Reply  Wed Jan 11 16:35:03 2017, gautam, Update, CDS, power glitch - recovery progress powerGlitchRecoveryJan2017.png
             Reply  Thu Jan 12 17:31:51 2017, gautam, Update, CDS, DC errors DCerror.png
                Reply  Fri Jan 13 21:41:23 2017, Koji, Update, CDS, DC errors 
Message ID: 12696     Entry time: Mon Jan 9 09:18:47 2017     In reply to: 0     Reply to this: 12699   12700
Author: Steve 
Type: Update 
Category: PEM 
Subject: power glitch 

There was a power glitch last night around 1:15am

The vacuum was not effected.

PSL laser turned on, PMC locked, PSL shutter opened and MC locked.

IR lasers at the ends turned on.

East arm air cond turned on.

The computers are all done.

The last power glitch was at Nov 3, 2016

 

 

Attachment 1: MondayMorning.png  571 kB  | Hide | Hide all
MondayMorning.png
ELOG V3.1.3-